A fly wheel has a speed of 300 r.p.m.

1 what is it angular velocity

2. What angle will be generated in 0.2 second

3 what much time is required for the wheel to rotate through 628 radians


1
Expert's answer
2022-05-16T05:16:01-0400

1.

"\omega=2\pi n\\
=2*3.14*300*1/60=31.4\:\rm rad/ s"

2.

"\theta=\omega t\\
=31.4*0.2=6.28\:\rm rad=360^\circ"

3.

"t=\theta/\omega\\
=628/31.4=20\:\rm s"